OpenStudy (anonymous):

Solve the following formula for x y = 8 mx -3 b x=

OpenStudy (littlebird):

It's either x=(y+3b)/8m Or x=(y+3)/8

OpenStudy (anonymous):

which one should I go with lol

OpenStudy (anonymous):

its y = 8x -3 y = 0 x= 3/8